ELECTRONIC TRANSITIONS OF THE HC_{4}S AND HC_{6}S RADICALS STUDIED BY LASER INDUCED FLUORESCENCE SPECTROSCOPY

\underline{MASAKAZU NAKAJIMA}, YOSHIHIRO SUMIYOSHI and YASUKI ENDO , Department of Basic Science, Graduate School of Arts and Sciences, The University of Tokyo, Tokyo 153-8902, Japan.

Laser induced fluorescence spectra of two new vibronic band systems have been observed by discharging a mixture gas, C_{2}H_{2} 0.5%/CS_{2} 0.3% diluted in Ar, in a supersonic jet. All the bands were assigned to ^{2}Π_{3/2} −^{2} Π_{3/2} transitions by measuring rotationally resolved spectra. Based on the chemical composition and the ground state combination differences, the spectral carriers were assigned to the HC_{4}S and HC_{6}S radicals. The homogeneous width of the excitation spectra of the HC_{6}S radical is larger than that of HC_{4}S by 0.02 cm^{−1} (FWHM), which was interpreted as the lifetime broadening of the excited state, corresponding to 270 ps. Vibronic origin of HC_{4}S is located at 500.5 nm, and that of HC_{6}S is at 589.6 nm. Main progressions of both the vibronic systems are assigned to the excitations of the C-S stretching modes, of which frequencies were determined to 505 cm^{−1} for HC_{4}S, 435 cm^{−1} for HC_{6}S.